L M N O P Q R S at 10 at A1 ✓ fx Part 2 C D E F G H I J K 7 The budget director for Bird House Unlimited, Inc., has gatheredA1 X fx Part 2 В. D E G H I J K L M N O P Q R S 31 32 33 34 35 Work in process-given in total of the 2 products Estimated invX V for Part 2 B C D E 2,100 units F G H $36 per unit I J K L M N O P Q R S T 43 Bird Feeder at 44 45 46 Direct Labor RequireFile Home Insert Draw Page Layout Formulas Data Review View Help Share Comments (i) PROTECTED VIEW Be careful—files from theV1 G H I J K L M N O P Q R S T ДА В D E F 18 19 Enter the appropriate cell references or formulas in the shaded cells below.A B F G H I J K L M N O P Q R S T Bird House Unlimited, Inc. Schedule to determine the required material for production For tG H I K L M N O P Q R S T AB c o E F Bird House Unlimited, Inc. Direct Labor Cost Budget For the Month Ending January 31, 202V1 X & fic с L M N O P Q R S T AB E F I K 98 99 6. Hint: Do not forget you have 2 products and 2 direct materials, so you wilV1 AB F G H I J K L M N O P Q R S T 119 120 D E Finished goods inventory, January 1, 2020 Cost of finished goods available fovi - X ✓ fx C D E F G H I K L M N O P Q R S T 1 ДА В 145 8. 146 147 148 149 150 151 152 Bird House Unlimited, Inc. Budgeted IApologies for the long question!
Answer #1

1.

Sales Budget
January
Birdhouse Birdfeeder Total
Unit Sales Volume 34000 25000 59000
Unit Selling Price $                            42 $                            70
Total Revenue from sales $             1,428,000 $             1,750,000 $                    3,178,000

2.

Production Budget
January
Birdhouse Birdfeeder Total
Expected Unit Sales 34000 25000 59000
Add : Desired Ending Finished Good Units 5300 2100
Total Required Units 39300 27100 66400
Less : Beginning Finished Good units 4900 2500
Required Production units 34400 24600 59000

3.

Direct Materials Budget
January
Wood Plastic Total
Required Units for production
Birdhouse 27520 =34400*0.8 17200 =34400*0.5 44720
Birdfeeder 29520 =24600*1.2 18450 =24600*0.75 47970
Add : Desired Ending inventory 3500 2800
Total Required Units 60540 38450 98990
Less : Beginning inventory 2600 3200
Total Units to be Purchased 57940 35250 93190
Unit Price $                        6.50 $                               0.90
Total costs for direct materials $                 376,610 $                          31,725 $    408,335

4.

Direct Labor Budget
January
Fabrication
Department
Assembly
Department
Total
Hours Required for Production
Birdhouse 8600 =34400*0.25 10320 =34400*0.3 18920
Birdfeeder 11070 =24600*0.45 8610 =24600*0.35 19680
Total hours 19670 18930 38600
Hourly Rate $                            14 $                                  10
Total cost of direct labor $                 275,380 $                        189,300 $    464,680

As per HOMEWORKLIB RULES we are supposed to answer 4 parts, i have answered 4, so kindly post other parts separately